OpenStudy (anonymous):

A particle with the mass of a proton is trapped in a box. The particle falls from the n=3 state to the n=2 state and emits 0.16 meV of energy. What is the size of the box? Please report your answer in nm.

OpenStudy (anonymous):

u need to calculate L^2=2.195*10^-66/3.245*10^-49 after sqrt(6.4*10^-18) and u got the answer :) an u need to convert the meters in nm(1m=10^9nm)